### v2

# Georgia Tech

Abstract
Multimodal Large Language Models (MLLMs) hold significant promise for revolutionizing traditional education and reducing teachers’ workload. However, accurately interpreting unconstrained STEM student handwritten solutions with intertwined mathematical formulas, diagrams, and textual reasoning poses a significant challenge due to the lack of authentic and domain-specific benchmarks. Additionally, current evaluation paradigms predominantly rely on the outcomes of downstream tasks (e.g., auto-grading), which often probe only a subset of the recognized content, thereby failing to capture the MLLMs’ understanding of complex handwritten logic as a whole. To bridge this gap, we release EDU-CIRCUIT-HW, a dataset consisting of 1,300+ authentic student handwritten solutions from a university-level STEM course. Utilizing the expert-verified verbatim transcriptions and grading reports of student solutions, we simultaneously evaluate various MLLMs’ upstream recognition fidelity and downstream auto-grading performance. Our evaluation uncovers an astonishing scale of latent failures within MLLM-recognized student handwritten content, highlighting the models’ insufficient reliability for auto-grading and other understanding-oriented applications in high-stakes educational settings. In solution, we present a case study demonstrating that leveraging identified error patterns to preemptively detect and rectify recognition errors, with only minimal human intervention (e.g., with 3.3% assignments routed to human graders while the rest to GPT-5.1 grader), can effectively enhance the robustness of the deployed AI-enabled grading system on unseen student solutions.
EDU-CIRCUIT-HW: Evaluating Multimodal Large Language Models on Real-World University-Level STEM Student Handwritten Solutions
Weiyu Sun Georgia Tech wsun355@gatech.edu          Liangliang Chen Georgia Tech liangliang.chen@gatech.edu          Yongnuo Cai Georgia Tech ycai350@gatech.edu
Huiru Xie Georgia Tech hxie77@gatech.edu          Yi Zeng Virginia Tech yizeng@vt.edu          Ying Zhang Georgia Tech yzhang@gatech.edu
1 Introduction
Figure 1: An exemplary STEM student handwritten solution auto-grading workflow: MLLM (e.g., Gemini-2.5-Pro) recognizes ①, ②, ③, and other visual contents of the solution, which serve as inputs for the consequent AI grading. Notably, recognition errors (① and ②) may not influence the grading if they fall outside the specific grading criteria. As a result, this successful grading masks underlying recognition failures, leading users to overestimate the MLLM’s visual understanding performance on the student solution.
Recent cutting-edge multimodal large language models (MLLMs) have demonstrated near-human performance over a wide range of low-stakes, daily visual understanding tasks. However, as the focus shifts to high-stakes, precision-critical domains, the visual understanding of these MLLMs remains insufficiently robust and inadequately evaluated. One particularly challenging and underexplored setting is university-level STEM (Science, Technology, Engineering, and Mathematics) student handwritten solution understanding, where the model must interpret the complex, interleaved “visual language” consisting of unconstrained handwriting, intricate mathematical derivations, and hand-drawn diagrams. The interpreted content further serves as input to consequent downstream tasks such as auto-grading (Kortemeyer et al., 2024; Khrulev, 2025) as shown in Figure 1. To this end, a reliable and trustworthy MLLM-based workflow to understand student handwritten solutions is a longstanding ambition in AI-enabled education.
Despite the potential of MLLMs, evaluating their effectiveness in real-world university settings faces two critical hurdles. First is the scarcity of proper authentic data. Most existing benchmarks fall short of capturing the complexity of real-world student handwritten work. For instance, many benchmarks (Xie et al., 2023; Gervais et al., 2025) focus on isolated visual elements (e.g., a single mathematical expression in Figure 1 ③) rather than the tightly interleaved reasoning processes characteristic of authentic student works (e.g., the entire input student solution in Figure 1). In addition, these benchmarks often omit or oversimplify hand-drawn diagrammatic understanding (e.g., ① and ② in Figure 1), despite diagrams playing a critical role in many STEM domains. Moreover, many prior benchmarks (Baral et al., 2025) target relatively low-difficulty settings such as K–12-level math. As a result, they fail to reflect the challenges posed by college STEM handwritten assignments. Consequently, such benchmarks provide limited insight into the true visual understanding and reasoning capabilities of MLLMs under realistic, high-stakes educational conditions.
Second, there is a fundamental misalignment in evaluation paradigms. Most existing benchmarks for student handwritten solution understanding focus on a single downstream objective (e.g., VQA (Baral et al., 2025) or auto-grading). While these task-specific evaluations are valuable, such downstream tasks often probe only a subset of the recognized content, leaving many recognition errors unobserved if they do not affect the specific task outcome. One such example is shown in Figure 1, where recognition errors in ① and ② are not reflected in the final grading report due to the rubric limitations. Crucially, these “negligible” errors in the auto-grading task may be catastrophic for other downstream tasks, such as circuit-to-netlist transformation (Xu et al., 2025).
To bridge the abovementioned gaps, we introduce EDU-CIRCUIT-HW, a diagnostic dataset grounded in real-world university STEM education, which comprises 1334 authentic student handwritten solutions to complex circuit analysis problems. To enable fine-grained MLLM visual recognition ability analysis, we curated near-verbatim transcriptions for a data subset and utilize them as the reference to capture different MLLMs’ errors in their recognized student solutions. Moreover, we build a taxonomy for these captured errors and evaluate their cascading impact on a downstream auto-grading task.
Our evaluation reveals a pervasive presence of latent recognition errors beneath current grading outcomes. Although these failures may remain dormant under coarse assessment criteria (e.g., binary correctness check), they become increasingly detrimental as grading granularity refines (e.g., specific, targeted score deduction), where higher precision inevitably exposes these underlying defects and degrades grading quality. This finding indicates that reliable student solution understanding and its applications remain a distant goal. In solution, we demonstrate the feasibility to detect and suppress unseen potential recognition failures via our identified error patterns in a case study.
Overall, our contributions are summarized as follows:
i) We release EDU-CIRCUIT-HW, a dataset of 1300+ authentic university-level STEM solutions, facilitating the evaluation of MLLMs’ visual understanding capabilities in real-world educational scenarios.
ii) We conduct a comprehensive evaluation of various cutting-edge MLLMs’ capabilities in handwriting understanding and downstream auto-grading on EDU-CIRCUIT-HW via a proposed automated diagnostic workflow. By establishing a fine-grained taxonomy of visual perception failures, we systematically analyze the cascading impact of recognition errors on the downstream auto-grading, which uncovers the latent risks concealed by seemingly robust downstream performance.
iii) We conduct a targeted case study to demonstrate that the identified error patterns can be leveraged to detect and prevent MLLMs’ recognition failures over unseen student solutions, thereby enhancing the reliability of deployed grading systems.
2 The EDU-CIRCUIT-HW Dataset
This section introduces the EDU-CIRCUIT-HW dataset for evaluating MLLMs’ capabilities on university-level STEM handwritten solution recognition and downstream auto-grading tasks.
2.1 Handwritten Solution Collection
The EDU-CIRCUIT-HW dataset consists of 1334 handwritten homework solutions from 29 students in an undergraduate-level circuit analysis course at a large, public, research-intensive institution in the Southeast United States during the Spring 2025 semester. The homework problems are all from the textbook (Svoboda and Dorf, 2013) with the topics varying from the basic circuit concepts and elements to advanced topics such as the first- and second-order circuit analyses. Each handwritten solution figure in this dataset corresponds to one student’s submission to a specific homework problem. In the pre-processing step, we removed the personally identifiable information of students, including students’ names and university IDs, from the solution images. All the students’ solutions are associated with expert-labeled reference grades in five different aspects, making the dataset able to be used to verify the performance of MLLM-enabled auto-graders, as illustrated in Section 3.2. Since students may leverage matrix theory, calculus, complex operations, and hand-drawn diagrams when solving the circuit problem, the benchmarks and analyses built on the EDU-CIRCUIT-HW can also be extended to other related STEM areas beyond circuit analyses.
2.2 Dataset Statistics
There are two data groups in the EDU-CIRCUIT-HW dataset: observation set and test set. The observation set consists of 513 handwritten solutions from 11 students. Each solution in the observation set is associated with detailed grades and image recognition results, of which both are verified and proofread manually by experts, as detailed in Section 3. Particularly, the expert-proofread recognition contains a near-verbatim transcription of all student-handwritten content, including natural-language descriptions for any non-textual elements such as hand-drawn circuit diagrams with annotations or function graphs. Since the recognized content is manually verified, the observation set can be regarded and utilized as a “training set” from which we can summarize the recognition patterns and use them to further improve the performance of both recognition and automated grading, as illustrated in Sections 35. In addition to the observation set, the 821 samples from the remaining 18 students constitute the test set, in which each handwritten solution is associated with a ground truth grade but not expert-verified recognition. Table 1 summarizes some key attributes of these two data groups.
Attribute Observation Set Test Set
Number of Unique Questions 62 62
Number of Students 11 18
Number of Samples 513 821
w/ Verified Recognition
w/ Ground Truth Grade
Student’s Solution Accuracy 72.12% 71.86%
Table 1: Key attributes of the observation and test sets in EDU-CIRCUIT-HW.
3 Handwritten Solution Recognition and Automated Grading Framework
Section 2 describes the collection and basic statistics of the EDU-CIRCUIT-HW dataset. In this section, we first detail how the handwritten solution can be effectively recognized by prompting MLLMs strategically. We then describe how the recognition errors can be classified into four well-defined categories and propose a method to investigate their impact on the downstream auto-grading task.
3.1 Handwritten Solution Recognition
Reliable handwritten solution recognition is essential and further affects the downstream task, such as homework auto-grading. However, even the most cutting-edge MLLMs perform poorly on STEM handwritten homework solution recognition tasks (Baral et al., 2025; Caraeni et al., 2024).
In this section, we analyze MLLMs’ recognition performance in a more fine-grained view. To this end, after obtaining MLLMs’ recognition results for the observation set, we build an LLM-as-a-judge pipeline to locate all potential recognition errors, which directly reflect the handwritten segments that the MLLM has trouble recognizing, as shown in Figure 2. Specifically, we first prompt Gemini-2.5-Pro to recognize the image-based students’ solution submissions and output the resulting textual descriptions in the Markdown format. In this process, the equations and texts are prompted to be recognized verbatim, and the diagrams, if existing, are interpreted in natural language that contains the full diagram information, such as the circuit topology and students’ annotations. The recognition results are then manually proofread against the original handwritten images by experts who are also responsible for rectifying the identified recognition errors, such as wrong diagram depictions or equations, by saving the rectified version in new Markdown files.
Figure 2: The demonstration on our recognition error detection. Given a handwritten solution, the Refer to caption MLLM’s recognition result is compared with the Refer to caption expert-verified transcriptions in a full-text level, after which all discrepant items within the recognition result are listed as recognition errors. Note that we regard two items as aligned when they are semantically equivalent in engineering education (e.g., Refer to caption KCL: “out” and Refer to caption KCL: @ “out” are regarded as semantically aligned despite the additional “@”).
The expert-proofread Markdown files serve as the ground truth to assist an LLM-judger, which we use Gemini-2.5-Pro, to capture the potential recognition errors from the evaluated MLLM. Specifically, for each handwritten solution, we provide the LLM-judger with both the verified recognition and the evaluated MLLM’s recognition and ask the LLM-judger to list all potential discrepant recognition items between them. Note that an item in this scenario can be either a sentence or an equation. Some item examples can be found in Figure 19 in Appendix D.
3.1.1 LLM-as-a-Judge Method Validation
In order to verify the LLM-judger’s reliability, we further manually check the recognition discrepancies and compare the identified discrepancies with those listed by the LLM-judger. The comparisons are made at two levels: sample level and item level. The sample level aims to measure the agreement between the LLM-judger and human experts in determining whether a handwritten recognition result contains any recognition errors. We use a binary indicator to describe this agreement. The item level focuses on whether each individual recognition error annotated by human experts (e.g., the discrepant case shown in Figure 2) can be detected by the LLM-as-a-judge method. At this level, we evaluate error detection performance using the precision, recall, and the F1 score, i.e.,
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (1)
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (2)
$\text{F1 Score}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (3)
where true positives (TP) are recognition errors identified by both the human expert and the LLM-judger, false positives (FP) are errors reported by the LLM-judger but not annotated by the human experts, and false negatives (FN) are errors annotated by the human experts but missed by the automated LLM-judger.
The comparison is conducted with over 186 recognized handwritten solution samples containing more than 5000 items. Specifically, we sample one student’s solution randomly for all the 62 problems in our dataset and prompt 3 commercial MLLMs (GPT-5.1, Claude-4.5-Sonnet, and Qwen3-VL-PLUS) to make recognitions. For each recognized result, an expert manually annotates all recognition errors (like those discrepant items in Figure 2). These annotations are then used to evaluate the performance of LLM-judger at both sample and item levels. The results are shown in Table 2, which indicates both high sample-level accuracy (larger than 0.95 on all three models) and item-level consistency (precision, recall, and the F1 score are all close to or larger than 0.9 on three models). A closer inspection also indicates that false positives and false negatives are predominantly associated with ambiguous items rather than reflecting systematic misjudgments. These results suggest that the LLM-as-a-judge pipeline is highly reliable and thus enables automated evaluation of handwritten recognition performance of different MLLMs with human-level consistency.
Model Sample Item-Level Metrics
Accuracy TP FP FN Precision Recall F1 Score
GPT-5.1 0.9516 97 6 10 0.9417 0.9065 0.9238
Claude-4.5-Sonnet 0.9839 137 13 15 0.9133 0.9013 0.9073
Qwen3-VL-Plus 0.9516 62 7 5 0.8986 0.9254 0.9118
Total 0.9624 296 26 30 0.9193 0.9080 0.9136
Table 2: Validation results of the automated recognition error detection via LLM-judgers against expert annotations across 186 handwritten solutions from three MLLMs.
3.1.2 Taxonomy on Recognition Errors
Recognition errors vary from trivial character mistakes to severe logical misunderstandings. To gain detailed insights, we further divide the recognition errors into four categories in Table 3. An LLM is prompted with clear category definitions to categorize all discrepant items in an automated way. With this taxonomy, we can observe the distribution of errors among different recognitions, which reveals the weaknesses of current MLLMs.
Category Introduction & Example
Symbolic & Character Misreading alphanumeric glyphs, operators, or units. Example: $20000\to 2a000$; $V\to V$
Structural & Notational Wrong mathematical layout or variable consistency. Example: $\text{R}=\frac{\frac{1}{8}}{\frac{1}{8}+\frac{1}{16}}\Omega\to\text{R}=\frac{8}{8+16}\Omega$
Diagrammatic Misinterpretation of diagram space attributes or annotations. Example: Erroneous circuit topology.
Textual & Logical Mislabeling context or sequence of reasoning. Example: Omission of critical derivation steps
Table 3: Taxonomy of recognition errors. More examples with image illustrations are available in Tables 811 in Appendix D.5.
3.2 Downstream Task: Automated Grading
Section 3.1 described the methods of handwritten solution recognition and detection for recognition errors. In this section, we will explore how the upstream recognition errors affect the downstream task, which is helpful to minimize the negative effect of the recognition errors and boost the downstream task performance.
In this work, we consider auto-grading as the downstream task, a process widely implemented across educational institutions worldwide. Specifically, we use an LLM-based grader, which is provided with MLLM-recognized student solutions and the problem context to assess student performance on the given problem. In our settings, the problem context includes the problem description, reference solution, and grading rubrics. The auto-grader’s reports are then compared with expert-verified reference grades available in the dataset, as described in Section 2.1.
In the EDU-CIRCUIT-HW dataset, each student’s handwritten solution was evaluated by a human expert based on a multi-dimensional rubric that covers five distinct perspectives of assessment as shown in Table 4. With these five grading aspects, the rubrics are defined to be problem-specific by using a unique rubric for each problem in our dataset. Given the rubric, the human expert will deduct scores for those students’ submissions violating the rubric per perspective. For example, “{E:0.02pts, U:0.01pts}” indicates that the student might use a wrong equation (e.g., incorrect Ohm’s Law $R=\frac{I}{V}$”) and an incompatible unit (e.g., using “voltage” instead of “ampere” to describe a current). These expert-verified grading results can serve as ground truth to evaluate the performance of the auto-grading pipeline.
Type Description
E Equation Error: Incorrect or inapplicable equations
M Method Error: Flawed problem-solving logic
U Unit Error: Missing or incorrect units
C Calculation Error: Wrong arithmetic with the correct equation
NC Incomplete Solution: Incomplete answer or missing steps
Table 4: The rubric for grading students’ work. A student’s solution can violate multiple standards in the table and get scores deducted separately. A rubric example can be found in Figure 21 in Appendix D.
Figure 3: The proposed auto-grading pipeline. The “vanilla grading pipeline” in green box is a widely-used auto-grading paradigm in AI-enabled education field. In our experiment (Section 4), we detect and analyze recognition errors within “MLLM Recognized Text” utilizing the automated pipeline introduced in Section 3.1, and further evaluate “Grading Report” by benchmarking it against the “Expert’s Grading Report” in yellow box. In pink box, we present an exemplary regrading module where we let the human experts summarize the detected recognition error patterns and leverage it ((detailed in Figure 25 and 26)) to filter out potential recognition errors within “MLLM Recognized Text”, and let LLM/human regraders to reassess those solutions flagged with containing potential recognition error, more details can be found in Section 5.
[…]
Graduate 83.63 82.46 81.29
GPT-5.1 Human Expert 89.47 78.36 74.46
Gemini-3-Preview 37.62 0.61 87.91 78.17 74.27 7.60%
Gemini-2.5-Pro 53.52 1.23 85.58 73.68 69.40 14.72%
GPT-5.1 71.54 2.05 77.78 65.50 61.99 17.89%
Qwen3-VL-Plus 61.72 1.38 80.90 68.62 65.11 16.67%
Claude-4.5-Sonnet 80.70 2.76 77.58 63.16 59.84 18.05%
Qwen3-VL-8B-Thinking 85.43 2.79 75.05 61.01 56.92 19.60%
Table 5: Overall evaluation results of different MLLMs’ recognition quality on student handwritten works in EDU-CIRCUIT-HW (the observation set), and their impact (i.e., Error Impact Rate (EIR)) on the downstream grading performance. We further add two baselines: (1) A Graduate student grades directly on all handwritten samples and (2) Human Expert recognizes and provides the “gold standard” recognized text for the LLM grader for reference. For metrics, we adopt Sample Error Rate (SER) and Average Error Count (AEC) to evaluate recognition quality, and binary, type, and point agreements to evaluate grading performance. The bold values are the best scores across different metrics. The detailed taxonomic EIR analysis over different error categories is detailed in Figure 4.
4 Experiments
This section shows the experiments on the observation set that includes the expert-proofread recognition results. Specifically, we conduct detailed analyses of recognition errors across various cutting-edge MLLMs and examine their relationship to downstream auto-grading performance.
4.1 Experiment Setup
In this experiment, we analyze the recognition capabilities of different cutting-edge MLLMs on students’ handwritten solutions in the observation set of EDU-CIRCUIT-HW. Specifically, we use the LLM-as-a-judge detector (introduced and validated in Section 3.1) to identify recognition errors in the “MLLM recognized text” as illustrated in Figure 3. For each model, we count the number of recognition errors, categorize their types, and examine their relationship with downstream grading outcomes. All grading results are obtained using the vanilla grading pipeline in Figure 3.
We evaluate five closed-source commercial models (Gemini-3-Pro-Preview, Gemini-2.5-Pro, Qwen3-VL-PLUS, Claude-4.5-Sonnet, and GPT-5.1) and one open-source model (Qwen3-VL- 8B-Thinking) as MLLM recognizers in Figure 3. Additionally, we include an oracle baseline in which expert-proofread transcriptions are provided to the LLM grader, representing grading performance under perfect recognition. We use GPT-5.1 as the LLM grader in all settings. More details can be found in Appendix D.
Comparison With Human Grader In addition, we include the grades assigned by the graduate teaching assistant (denoted as Graduate in Table 5) in the data collection course. This serves as a strong baseline, enabling us to assess how far current cutting-edge MLLM-based auto-graders have progressed in handling relatively challenging university-level problems.
Figure 4: Comparisons of different MLLMs’ recognition error counts over 4 error categories (defined in Table 3) and the corresponding error impact rate (EIR) to the downstream LLM grading performance in Table 5. The annotation above each bar presents the error count (line 1) and the EIR (line 2 in gray) of this model in the category.
Recognition Quality and Grading Metrics We evaluate recognition performance at two granularities: Sample Error Rate (SER) and Average Error Count (AEC). SER measures the proportion of recognized texts containing at least one recognition error, while AEC represents the item-level average number of errors per handwritten solution.
To evaluate grading performance, based on the rubric introduced in Table 4, we measure the agreement between the LLM and expert grading reports at three levels: (1) Binary Agreement: Evaluates whether the model correctly identifies the presence of any student mistakes; (2) Type Agreement: Requires consistency in the flagged error types (E, M, U, C, NC); (3) Point Agreement: The strictest metric that requires exact matches in both error types and their corresponding point deductions. For example, given an LLM grading report {E:-0.1pts,M:-0.3pts} and an expert grading report {E:-0.1pts,M:-0.2pts}, binary agreement is satisfied since both identify errors, and type agreement is also satisfied since both flag the same error types (“E” and “M”). However, point agreement is not satisfied due to the discrepancy in the deducted score for perspective “M”.
Error Impact Analysis To further investigate the propagation of errors from visual recognition to the final grading, we introduce the Error Impact Rate (EIR). This metric quantifies the proportion of item-level recognition errors that directly lead to a downstream grading discrepancy. It is calculated as the number of recognition errors that result in a grading error divided by the total number of item-level recognition errors identified.
4.2 Results and Findings
The experimental results in Table 5 reveal several key insights. First, the findings empirically support our argument that task-centric evaluation alone fails to expose the full spectrum of visual recognition errors. Although automated grading is a multi-faceted downstream task, it inherently masks a significant portion of recognition failures. For instance, while Gemini-3-Preview exhibits a Sample Error Rate (SER) of 37.62%, its Error Impact Rate (EIR) is merely 7.60%. Nevertheless, the recognition errors have amplified negative impacts on the downstream fine-grained auto-grading tasks. According to Table 5, with SER and EIR increasing from Gemini-3-Preview to Qwen3-VL-8B-Thinking, the grading performance gaps increase from 12.86% in the binary agreement to 17.16% and 17.35% in the type and point agreements which require the correct identification of error types and precise point deductions. Additionally, as the overall recognition deteriorates, the EIR increases from 7.60% to 19.60%. These trends imply that stronger visual understanding captures more contextually significant information from student work, thereby reducing the likelihood of catastrophic error propagation.
Second, when comparing LLM-based graders with the human graduate grader, we observe a clear divergence between coarse-grained and fine-grained assessments. State-of-the-art MLLMs (e.g., the Gemini series) outperform the graduate grader in the binary agreement. However, this advantage diminishes when facing more strict criteria. On the type and point agreements, even the strongest MLLMs lag behind the human grader with non-negligible performance gaps. This indicates that current MLLMs still struggle to reliably capture and reason over critical, fine-grained details in handwritten solutions, highlighting a persistent challenge in deploying LLM-based systems for precision-critical educational evaluation.
Recognition Error Taxonomy Based on the taxonomy defined in Table 3, we categorize the identified recognition errors to analyze their distribution and impact, as summarized in Figure 4. Unsurprisingly, recognition errors are most prevalent at the surface level, particularly in the Symbolic & Character category, while more complex Diagrammatic and Textual & Logical errors are less frequent. Beyond this, we observe that different MLLMs exhibit heterogeneous visual recognition capabilities. For instance, despite its smaller scale, Qwen3-VL-8B-Thinking demonstrates competitive performance in the Diagrammatic category (98 errors), even surpassing larger models like Gemini-2.5-Pro (103 errors). Conversely, commercial models generally excel in the Textual & Logical category, showcasing their superior ability to interpret high-level semantic relations and textual explanations.
In addition, the Error Impact Rates (EIRs) in Figure 4 reveal the sensitivity of grading to different error types. Notably, surface-level errors in Symbolic & Character have a more direct impact on grading outcomes, with EIRs consistently around 20%. In contrast, identification failures in more complex Diagrammatic and Textual & Logical categories generally exhibit lower EIRs, frequently falling below 10%. This suggests that current auto-grading pipelines remain heavily reliant on symbolic matching rather than structural or diagrammatic reasoning. However, while diagrammatic errors appear less “impactful” for simple grading, they remain catastrophic for downstream STEM-specific tasks such as circuit-to-netlist generation (Rachala and Panicker, 2022) or complex problem solving (Chen et al., 2025; Skelic et al., 2025). These findings underscore that achieving truly robust and comprehensive STEM handwriting understanding still remains a significant challenge.
5 Case Study
To assess the practical utility of our recognition error analysis, we conduct a deployment-oriented case study that integrates our empirical recognition error diagnostics into a realistic human-in-the-loop grading workflow (i.e., the “regrading module” in Figure 3). Specifically, we first collect identified recognition errors from the observation set and summarize their patterns. These patterns are then used to prompt the LLM to identify potential errors within the MLLM-recognized student solutions in the test set, which simulate unseen scenarios in practice. Meanwhile, considering that the distinction between genuine recognition errors and students’ own mistakes can sometimes be ambiguous (due to handwriting or logical issues), we further ask the LLM to provide a confidence level (“high” or “low”) for each identified error. Solutions flagged with low-confidence errors are routed directly to the course teaching assistant for manual grading, while the remaining solutions are re-graded by the LLM grader given the detector’s error report. More details can be found in Appendix D.3.
Workflow Visual Recognizer Grading Perf. (%) $\uparrow$ Regrading Rate (%)
Binary Type Point LLM Human
[…]
Table 6: Performance and regrading statistics before and after applying the regrading module for student solutions in the test set. LLM and Human columns denote the percentage of solutions assigned to LLM and teaching assistant for regrading, respectively.
Deployment Setup We use the “vanilla grading pipeline” in Figure 3 for the first round student solution grading, where GPT-5.1 and Gemini-2.5-Pro are tested as the visual recognizer separately. Subsequently, we apply the “regrading module” to regrade all solutions that received point deductions in the first round. This design is motivated by both error characteristics and deployment costs: recognition errors primarily result in false positive grading penalties, and cases with point deductions generally constitute a small fraction of student solutions. The LLM grader, LLM re-grader, and recognition error detector are all implemented using GPT-5.1.
Performance Gain from Regrading Table 6 indicates that the “regrading module” significantly improves grading performance, particularly in type and point agreement, which necessitate precise alignment of fine-grained details. Notably, with fewer than 5% solutions requiring human intervention, the regrading achieves performance comparable to the setting where human experts serve as visual recognizers for all solutions in Table 5. This demonstrates that despite the low error impact ratio, targeted recognition error suppression remains indispensable for achieving expert-level accuracy in complex STEM evaluation.
6 Conclusions
In this work, we introduce EDU-CIRCUIT-HW, a real-world benchmark comprising over 1,300 authentic university-level STEM handwritten student solutions. Leveraging expert-verified verbatim handwritten solution transcriptions and grading reports, we comprehensively evaluate the visual understanding capability of MLLMs and their cascading impact on downstream auto-grading. The evaluation uncovers a substantial volume of latent recognition errors that may be exposed when rubrics become stricter and more professional, highlighting the significant challenges in deploying reliable MLLM-based educational applications. Furthermore, in response to these challenges, we propose and validate a human-in-the-loop grading workflow to mitigate the impact of recognition failures, which significantly enhances system robustness with minimal human intervention. We hope our work can serve as a cornerstone for developing more trustworthy AI-enabled educational technologies.
7 Data Scope: Beyond Circuit Analysis
As introduced in Section 2.1, the student handwritten data in EDU-CIRCUIT-HW was collected from a circuit-analysis-related university-level course. However, we argue that this does not imply that our data merely serves as a proxy for a narrow subset of STEM challenges. In effect, solving the problems included in our dataset requires interleaved reasoning across multiple foundational disciplines, including calculus, differential equations, and physics. These subjects are formal prerequisites for the course at our institution, meaning that students must demonstrate competence in them before enrolling in the target course. In Figure 10 and 11, we provide two recognized handwritten solutions corresponding to Problems 9.8-2 and 10.6-4 in the textbook by Svoboda and Dorf Svoboda and Dorf (2013). These two examples indicate that our EDU-CIRCUIT-HW dataset is also relevant to foundational STEM subjects such as physics, differential equations, calculus, and linear algebra. Due to copyright restrictions, we do not provide the original problem statements; instead, we briefly summarize them as follows:
i) The first example is a recognized handwritten solution to Problem 9.8-2 in the textbook [1]. The circuit diagram includes two resistors, one capacitor, one inductor, and an independent voltage source with voltage 2u(t)12u(t)-1. The task is to determine the current $i(t)$ through the inductor for $t>0$. The student’s solution involves fundamental circuit laws, including Kirchhoff’s Current Law (KCL) and Kirchhoff’s Voltage Law (KVL), the formulation and solution of a second-order linear differential equation, derivation of the characteristic equation in the Laplace domain, and representation of the final current as the sum of two exponential terms and one constant term. These elements indicate that the benchmarking results obtained from our dataset are not limited to circuit analysis content alone.
ii) The second example is a recognized handwritten solution to Problem 10.6-4 in the textbook [1]. The circuit diagram contains an independent voltage source represented by the phasor $48\angle 75^{\circ}$ and five impedance elements with impedances $40+j15\Omega$, $25-j50\Omega$, $32+j16\Omega$, $j50\Omega$, and $j40\Omega$. The task is to determine the values of three phasor mesh currents. The student’s solution involves three mesh-current equations forming a system of linear equations. When expressed in matrix form, the system includes complex-valued coefficients. Recognition and evaluation of this solution therefore require not only circuit analysis knowledge but also competence in linear algebra.
We hope the above clarification of our data scope can help the following researchers better understand and utilize our EDU-CIRCUIT-HW dataset.
8 Limitations
This work focuses on the analysis on handwritten solution visual understanding capability in a university-level circuit analysis course. Although such a course represents a comprehensive STEM subject that integrates concepts from mathematics, physics, and engineering, the problem set and accompanying diagrams are mainly centered on circuit-related analysis. Other diagram types strongly associated with abstract mathematics (e.g., complex geometric graphs) are underrepresented in our dataset. Therefore, the conclusions drawn in this study may not fully generalize to handwritten recognition tasks involving substantially different diagram modalities.
In addition, in our analysis of how upstream recognition errors propagate to downstream tasks, we mainly consider auto-grading as the primary downstream application. Although auto-grading is a practical and representative task in educational settings, we do not investigate the impact of recognition errors on other downstream tasks, such as visual question answering (VQA), which may exhibit different sensitivity patterns to recognition inaccuracies.
Finally, the grading labels and rubrics in our dataset are provided by few doctoral experts with extensive familiarity with the course content. However, human evaluation and rubric-crafting in open-ended STEM problem-solving are inherently subjective to some extent. Despite our efforts to ensure consistency through detailed grading criteria and rigorous verification grounded in mathematical logic and domain expertise, potential subtle subjective biases may still exist in the grading scores.
Overall, we view these limitations as opportunities for future work to extend the analysis to broader subject domains, additional downstream tasks, and alternative assessment protocols.
9 Ethics Considerations
Data Collection Prior to data collection, all participating students signed informed consent forms, authorizing the collection and usage of their homework solutions and grading records for research and publication purposes. All data were anonymized to remove Personally Identifiable Information (PII) of students (e.g., we identify students in our dataset via “student + index” (student 1, student 6, etc) instead of their true names for privacy protection.), and the collection was approved by the Institutional Review Board (IRB).
Copyright and Licensing The 62 questions used in this study are sourced from the textbook Svoboda and Dorf (2013), with their corresponding IDs listed in Table 7. To comply with copyright regulations, EDU-CIRCUIT-HW does not distribute and publicize the original problem statements, figures, or image-based hints. Researchers who wish to replicate our experiments involving the original question content are advised to consult the textbook themselves using the provided IDs. However, for all questions listed in Table 7, we have independently developed the reference solutions and comprehensive grading rubrics (as exemplified in Figure 21). These annotations are original contributions and are fully included in EDU-CIRCUIT-HW.
Reliability Risks in Educational Settings As analyzed in our work, MLLMs adopted in current auto-grading systems present unreliable recognition qualities on university-level STEM student solutions, which could propagate to downstream grading performance and thereby lead to unfair or incorrect assessments if deployed without sufficient human oversight. We emphasize that our dataset and analysis are intended for diagnostic and research purposes, and that any practical deployment should include human-in-the-loop verification and error detection mechanisms, as shown in the case study.
Appendix A Related Work
A.1 Handwriting Text Recognition
Despite the prevalence of MLLMs, they still face several challenges. One notable issue is the difficulty in accurately recognizing students’ handwriting and converting it into clean, machine-readable text. Although Optical Character Recognition (OCR) Hamad and Kaya (2016) and visual understanding techniques can reliably process standard text and simple, one-line mathematical expressions Wu et al. (2022), their performance degrades when confronted with complex deductions, hand-drawn auxiliary plots, and uncommon symbol combinations. This leads to low confidence in the automated evaluation of student work Yang et al. (2025); Caraeni et al. (2024), thereby increasing the need for manual verification Yang et al. (2025). Consequently, it is meaningful to analyze in detail the impact of recognition errors on auto-grading systems and to develop targeted solutions for these shortcomings.
A.2 Handwriting Text Recognition Correction
Handwriting Text Recognition (HTR) correction plays a crucial role in identifying and rectifying recognition errors, thereby improving the overall performance of the recognition pipeline. For example, Pavlopoulos et al. (2023) demonstrated that a sequential “detect-then-rectify” approach could enhance Greek text recognition accuracy to 97%. Similarly, Chen et al. (2023) employed an external LLM to automatically correct mathematical formulas by leveraging LaTeX grammar and semantic context, achieving an approximate 6% improvement in expression recognition accuracy on the CROHME dataset Mouchère et al. (2014).
Methods for error detection vary, including BERT-based models Peng et al. (2021) and confidence-based approaches Gupta et al. (2021). However, their effectiveness remains underexplored in the context of complex student handwriting, especially when accounting for practical constraints such as minimizing API call frequency and ensuring high-recall error detection (i.e., identifying as many errors as possible). This paper introduces a specialized error detection and correction pipeline tailored to student handwriting, developed and evaluated within the context of a university-level circuit analysis course, as detailed in the case study in Section 5.
A.3 Automated Grading
Recent auto-grading agents primarily leverage the capabilities of foundation models such as Gemini and GPT. For handwriting input, some researchers have explored end-to-end grading approaches that process handwritten submissions directly Caraeni et al. (2024); Khrulev (2025); Nath et al. (2025). However, the opaque nature and relatively lower performance of these models render them unreliable at the current stage. Alternatively, a more interpretable approach involves first recognizing the student’s handwritten solution and converting it into a structured text format (e.g., LaTeX or Markdown), which is then used as input for the grading model along with a prompt and predefined rubric. For instance, Liu et al. (2024) utilized GPT-4 to automatically grade a university-level mock mathematics exam. Kortemeyer et al. (2024) combined GPT-4V and Mathpix to evaluate student performance on a high-stakes thermodynamics exam. Similarly, Parsaeifard et al. (2025) developed a pipeline to automatically grade economics-related mathematics courses at a Swiss distance university. Yang et al. (2025) introduced an online auto-grading platform [^1] See https://www.pensieve.co/schools. that supports various academic institutions across subjects such as Computer Science, Mathematics, and Chemistry.
Prompt for categorizing the recognition errors based on the error taxonomy
Figure 5: Prompt used to categorize recognition errors according to the taxonomy defined in Table 3, where the four categories correspond to “Symbol & Character”, “Structural & Notational”, “Diagrammatic”, and “Textual & Logical” errors.
Appendix B Dataset Details and Ethics Statements
The EDU-CIRCUIT-HW dataset consists of the following components: (1) 1334 authentic student handwritten solutions collected from a university-level STEM course during the Spring 2025 term; (2) 62 reference answers and corresponding grading rubrics; (3) 513 expert-crafted, near-verbatim transcriptions of student solutions within the observation set; and (4) detailed records of recognition errors, categorized according to the taxonomy presented in Table 3. Further descriptions of these components are provided below.
(1) Student Handwritten Solution Collection. As summarized in Table 1, data were collected from 29 students enrolled in the course. Prior to data collection, all participants signed an informed consent form permitting the use of their homework submissions and grading scores for research purposes. All data were anonymized to remove Personally Identifiable Information (PII), and the collection process was approved by the Institutional Review Board (IRB). Handwritten assignments and corresponding grading reports were gathered throughout the Spring 2025 term. Notably, some students occasionally submitted assignments in LaTeX or other digital formats; these were excluded from the dataset. As a result, not all students in EDU-CIRCUIT-HW have handwritten solutions for every question.
(2) Questions, Reference Solutions, and Rubrics. The IDs of all 62 questions, sourced from the textbook Svoboda and Dorf (2013), are listed in Table 7. Due to copyright restrictions, EDU-CIRCUIT-HW does not include the original problem statements, figures, or image-based hints. Researchers wishing to replicate experiments involving the original content are encouraged to consult the textbook using the provided question IDs. For all questions listed in Table 7, we developed corresponding reference answers and comprehensive grading rubrics, both of which are included in EDU-CIRCUIT-HW. An example rubric is shown in Figure 21.
(3) Expert-Rectified Student Solution Recognitions. This component is described in greater detail in Appendix C and was previously introduced in Section 3.1.
(4) Recognition Errors and Taxonomy. As detailed in Section 3.1 and Appendix C, human experts identified recognition errors in the transcriptions produced by Gemini-2.5-Pro on the observation set within EDU-CIRCUIT-HW. These errors were then categorized using the taxonomy defined in Table 3, with classification performed by prompting Gemini-2.5-Pro using the prompt shown in Figure 5. In addition, we will release the recognition errors identified across all evaluated models (as shown in Figure 4), flagged by the LLM-based error detector and organized according to the same taxonomy, to support further research.
Chapter Index Problem Indexes
1 P1.5-2, P1.5-3
2 P2.4-6, P2.4-7, P2.5-1, P2.5-2
3 P3.2-1, P3.2-2, P3.2-6, P3.4-4, P3.6-1
4 P4.2-2, P4.2-5, P4.3-3, P4.4-1, P4.4-6, P4.5-2, P4.6-2, P4.7-1, P4.7-2, P4.7-6
6 P6.3-2, P6.3-3, P6.4-1, P6.4-2, P6.4-4, P6.4-5, P6.4-6
7 P7.2-2, P7.2-4, P7.3-2, P7.4-1, P7.4-3, P7.5-4, P7.5-5, P7.6-2, P7.7-2, P7.8-2
8 P8.3-1, P8.3-2, P8.3-4, P8.3-5, P8.3-9, P8.3-10, P8.4-1, P8.7-3, P8.7-5
9 P9.2-1, P9.3-3, P9.4-1, P9.4-2, P9.5-1, P9.6-1, P9.7-2
10 P10.3-2, P10.3-3, P10.4-1, P10.4-6, P10.6-1, P10.6-4, P10.7-1, P10.10-4
Table 7: Index of problems from Svoboda and Dorf (2013) corresponding to the data included in EDU-CIRCUIT-HW.
Appendix C Expert Handwriting Recognition
This section outlines the procedure for generating expert-verified handwritten recognition results, referred to as the “gold standard”, for student solutions. For all data in the EDU-CIRCUIT-HW observation set, we initially employed Gemini-2.5-Pro to recognize and transcribe the handwritten submissions into Markdown format, using the prompt shown in Figure 16. Subsequently, human experts carefully reviewed each transcription against the original handwritten images. Any recognition errors identified during this verification process were manually corrected by the experts. The resulting set of transcriptions serves as the gold standard and is assumed to be error-free for the experiments reported in Table 5.
To better illustrate the expert rectification process, we provide a representative example. Figure 6 presents the original student handwritten submission; Figure 12 shows the initial transcription generated by Gemini-2.5-Pro; and Figure 13 displays the expert-rectified version based on the model’s output. In these figures, recognition errors in the original transcription are highlighted in red, while the corresponding corrections in the expert-verified version are marked in green.
Figure 6: Screenshot of a handwritten solution from Student 16 for question P9.4-2, included in the observation set.
Figure 7: Screenshot of a student handwritten solution for question P8.3-10 from the test set. This example is used as the one-shot demonstration in the LLM-based recognition error detector described in Section 3.1 for evaluating the observation set.
Appendix D Experiment Details
D.1 Automated LLM-Enabled Recognition Error Detector
This section describes the implementation of the LLM-enabled automated recognition error detector, as introduced in Section 3.1. Leveraging the expert-verified transcriptions as the “gold standard”, we use Gemini-2.5-Pro to detect discrepancies between the recognition outputs produced by various MLLMs and the corresponding reference texts. The prompt used for this comparison process is illustrated in Figure 18, and consists of the following components:
  • target_content: The Markdown transcription generated by the target MLLM for a specific student handwritten solution.
  • label_content: The expert-verified Markdown transcription of the same student solution, generated through the pipeline described in Section C. This serves as the gold-standard reference for evaluating recognition accuracy.
  • examples[‘target’]: A sample transcription generated by GPT-5.1 for the student submission shown in Figure 7, with the transcribed output illustrated in Figure 14. To ensure fair evaluation (e.g., in Tables 2 and 5) and prevent data leakage, this example, corresponding to question P8.3-10 from Student 19, was selected from the test set, which lies outside the EDU-CIRCUIT-HW observation set.
  • examples[‘gt’]: The expert-rectified transcription corresponding to the student submission shown in Figure 7, as illustrated in Figure 15.
  • examples[‘result’]: A list of sample recognition errors and their corresponding corrections, derived by comparing examples[‘target’] with examples[‘gt’]. This list is shown in Figure 19, and the detector’s output follows the same structure.
  • examples[‘evidence’]: Detailed justifications for each recognition error identified in examples[‘target’], based on comparisons with examples[‘gt’]. These justifications are illustrated in Figure 20.
This prompt utilizes a one-shot strategy based on a representative example from the test set. Specifically, we intentionally paired a transcription produced by GPT-5.1 with an expert-rectified ground truth originally derived from Gemini-2.5-Pro. By incorporating outputs from different MLLMs in the one-shot example, the detector is exposed to the natural variability in how models represent the same visual content, for instance, through differing LaTeX notations for equivalent formulas or alternative phrasings for diagrammatic descriptions. This cross-model pairing enhances the robustness of the LLM-based detector by helping it distinguish between benign structural or stylistic differences and true recognition errors.
D.2 Vanilla Grading Pipeline
The “vanilla grading pipeline”, illustrated in Figure 3, comprises two main stages: (1) transcription of the student handwritten solution by an MLLM, such as Gemini-2.5-Pro, and (2) automated evaluation by an LLM-based grader. In the transcription stage, we utilize different MLLMs to convert handwritten student responses into structured text, guided by prompts specific to each model series (see Figures 16 and 17).
In the second stage, the LLM grader (i.e., GPT-5.1), configured with a “high” reasoning effort setting, evaluates the transcribed content through a two-round dialogue process. In the first round, the grader employs the prompt shown in Figure 22 to assess the submission based on four question-specific attributes: problem_statement, images_hint, final_answer (the reference solution), and rubric_snippet. Within EDU-CIRCUIT-HW, we provide the latter two components, developed by domain experts, while the problem_statement and images_hint can be retrieved from the textbook (Svoboda and Dorf, 2013) using the corresponding question indexes. The student_markdown field contains the transcribed student solution, which may originate from an MLLM (e.g., Figure 14) or a human expert, depending on the experimental setting. After the first round, the grader outputs an overall score deduction report (e.g., {E: -0.04 pts, C: -0.01 pts}).
In the second round, the LLM grader is prompted (as shown in Figure 23) to provide detailed justifications for the score deductions identified in the first round. The input to this stage includes the raw_scoring_json (e.g., {E: -0.04 pts, C: -0.01 pts}) generated earlier. This round concludes with the generation of a final grading report. A representative example is shown in Figure 9, which presents the LLM grader’s feedback on the student submission in Figure 8, evaluated using the rubric provided in Figure 21.
Figure 8: Screenshot of a student handwritten solution for question P9.7-2. A grading report example for this submission, generated by the LLM grader, is presented in Figure 9, based on the rubric shown in Figure 21.
A grading report example
Figure 9: Grading report for a student submission on question P9.7-2. The report includes score deductions in the categories of “Equation” (E) and “Calculation” (C), along with detailed explanations for each deduction and the corresponding evidence (e.g., evidence: \$\$B = 350\$\$) from the transcribed student solution.
Image Error Details
[Uncaptioned image] KVL: $5+4i_{1}\text{ ...}$
$\downarrow$
KVL: $15+4i_{1}\text{ ...}$
[Uncaptioned image] $e^{j89.37^{\circ}}$
$\downarrow$
$e^{j84.37^{\circ}}$
Table 8: Examples of the recognition errors in the “Symbolic & Character” category in Table 3. The incorrect recognitions are marked in red and the rectified versions are marked in green.
D.3 Regrading Module
This section provides additional details on the “regrading module” in Figure 3, which was implemented as part of the case study described in Section 5.
Image Error Details
[Uncaptioned image] $\frac{1}{\frac{1}{8~\text{mH}}+\frac{1}{8~\text{mH}}}=4~\text{mH}$
$\downarrow$
$\frac{1}{8~\text{mH}+8~\text{mH}}=4~\text{mH}$
[Uncaptioned image] $v_{c}=(1+\frac{R_{6}}{R_{4}})v_{b}$
$\downarrow$
$v_{c}=(\frac{1+R_{6}}{R_{4}})v_{b}$
Table 9: Examples of the recognition errors in the “Structural & Notational” category in Table 3. The incorrect recognitions are marked in red and the rectified versions are marked in green.
Image Error Details
[…]
$\quad\quad\quad\quad\downarrow$
[…]
[Uncaptioned image] The curve has a minimum between tt=0 and tt=1.
$\quad\quad\quad\quad\downarrow$
The curve has a minimum between tt=1 and tt=2.
[Uncaptioned image] The student circles two 4mH inductors (top right & rightmost) in red.
$\quad\quad\quad\quad\downarrow$
The student circles the rightmost 4mH inductor in red.
Table 10: Examples of the recognition errors in the “Diagrammatic” category in Table 3. The incorrect recognitions are marked in red and the rectified versions are marked in green.
Image Error Details
[Uncaptioned image] Node voltage analysis $\rightarrow$ KCL
$\downarrow$
Node voltage analysis $\rightarrow$ KVL
[Uncaptioned image] $i_{1}+i_{2}=\text{A}i_{1}$.
$1.333-4=\text{A}(-1.333)$
$\downarrow$
[Missing Equation].
$1.333-4=\text{A}(-1.333)$
Table 11: Examples of the recognition errors in the “Textual & Logical” category in Table 3. The incorrecrt recognitions are marked in red and the rectified versions are marked in green.
To implement this module, we first aggregated all recognition errors flagged by human experts within the observation set; the process for collecting these errors is detailed in Section C. Leveraging this prior knowledge, experts distilled a set of empirical criteria, presented in Figures 25 and 26, to guide the detection of potential recognition errors in the test set. These criteria were then integrated into the recognition error detector, which identifies likely transcription failures in student responses that received score deductions in the initial grading stage. For the case study, we use GPT-5.1 as the detector. Based on its output, the LLM regrader re-evaluates the affected student work using the prompt shown in Figure 24. During this process, the regrader is explicitly instructed to disregard any content flagged as a recognition error, even if it introduces discrepancies with the reference solution.
In cases where the detector is unable to definitively distinguish between a recognition error and a genuine student mistake (i.e., segments marked as “[uncertain]” in Figure 25), the corresponding solution is referred to a human grader for final evaluation. In our case study, this human grader is the doctoral teaching assistant for the course from which EDU-CIRCUIT-HW was derived.
D.4 Human and MLLM API Settings
Experts and Human Graders The graduate grader referenced in Table 5 is a graduate student who previously completed the course associated with EDU-CIRCUIT-HW and has also served as a teaching assistant for the class. All experts and human graders involved in the student work regrading are doctoral students with comprehensive mastery and deep subject-matter expertise in the course content.
Model Configurations All MLLMs used in our visual recognition experiments are accessed exclusively through their official API interfaces. To ensure reproducibility and reduce output variability, we set the temperature parameter to 0 for all models, except for GPT-5.1, whose temperature is fixed at 1 by the official setting222. Unless otherwise specified, all other hyperparameters (e.g., reasoning effort for GPT-5.1) remain at their default API settings. Given the low-temperature configurations, which result in highly stable outputs, and the practical constraints of deployment, where inference is typically conducted only once to control costs, we report experimental results based on a single run per model.
D.5 Examples about Recognition Error Taxonomy
In this section, we present several examples of recognition errors, categorized according to the taxonomy defined in Table 3. The examples are organized as follows:
  • Symbolic & Character: Representative recognition errors in this category are provided in Table 8.
  • Structural & Notational: Examples of errors involving equation structure or notation are shown in Table 9.
  • Diagrammatic: Recognition errors related to diagrams are illustrated in Table 10.
  • Textual & Logical: Examples of misinterpretations involving textual descriptions or logical reasoning appear in Table 11.
Example 1 for data scope clarification (A recognized handwritten solution to Problem 9.8-2; involving foundations from physics, differential equations, and calculus) For $t<0$: $i(0)=\frac{-1}{1+4}=-0.2\text{ A}$ $v(0)=\frac{4}{1+4}(-1)=-0.8\text{ V}$ For $t>0$: KCL $\rightarrow$ $\frac{v-V_{s}}{R_{1}}+C\frac{dv}{dt}+i=0$ KVL to right half $\rightarrow$ $v=R_{2}i+L\frac{di}{dt}$ plug KVL eq. to KCL eq.: $\frac{R_{2}i+L\frac{di}{dt}-V_{s}}{R_{1}}+C\left[R_{2}\frac{di}{dt}+L\frac{d^{2}i}{dt^{2}}\right]+i=0$ $\Rightarrow\frac{d^{2}i}{dt^{2}}+\frac{(L+R_{1}R_{2}C)}{R_{1}LC}\frac{di}{dt}+\frac{(R_{1}+R_{2})}{R_{1}LC}i=\frac{V_{s}}{R_{1}LC}$ $\Rightarrow\frac{d^{2}i}{dt^{2}}+5\frac{di}{dt}+5i=1$ Forced response is constant, $i_{f}=B$ so $1=\frac{d^{2}B}{dt^{2}}+5\frac{dB}{dt}+5B$ Therefore $B=0.2\text{ A}$ Characteristic eq: $s^{2}+5s+5=0$ Roots $\Rightarrow s_{1}=-3.62$, $s_{2}=-1.38$ $i_{n}=A_{1}e^{-3.62t}+A_{2}e^{-1.38t}$ $i(t)=A_{1}e^{-3.62t}+A_{2}e^{-1.38t}+0.2$ $v(t)=\left(4i(t)+4\frac{di(t)}{dt}\right)=-10.48A_{1}e^{-3.62t}-1.52A_{2}e^{-1.38t}+0.8$ At $t=0^{+}$, $0.2=i(0^{+})=A_{1}+A_{2}+0.2\Rightarrow A_{1}+A_{2}=-0.4$ $0.8=v(0^{+})=-10.48A_{1}-1.52A_{2}+0.8\Rightarrow-10.48A_{1}-1.52A_{2}=-1.6$ Using matrix calc: $A_{1}=0.246$ and $A_{2}=-0.646$ $i(t)=0.246e^{-3.62t}-0.646e^{-1.38t}+0.2\text{ A}$
Figure 10: An example in our dataset including a student’s solution on using foundations from physics, differential equations, and calculus to solve the target current.
Example 2 for data scope clarification (A recognized handwritten solution to Problem 10.6-4) Mesh 1: $(40+j15)i_{1}+(25-j50)(i_{1}-i_{3})-48\angle 75^{\circ}=0$ $(65-j35)i_{1}-(25-j50)i_{3}=48\angle 75^{\circ}$ Mesh 2: $48\angle 75^{\circ}+(-j50)(i_{2}-i_{3})+(32+j16)i_{2}=0$ $(32-j34)i_{2}+j50i_{3}=-48\angle 75^{\circ}$ Mesh 3: $j40i_{3}-(-j50)(i_{2}-i_{3})-(25-j50)(i_{1}-i_{3})=0$ $(-25+j50)i_{1}+j50i_{2}+(25-j160)i_{3}=0$ $\begin{bmatrix}-25+j50&j50&25-j160\\ 0&32-j34&j50\\ 65-j35&0&-25+j50\end{bmatrix}\begin{bmatrix}i_{1}\\ i_{2}\\ i_{3}\end{bmatrix}=\begin{bmatrix}0\\ -48\angle 75^{\circ}\\ 48\angle 75^{\circ}\end{bmatrix}$ $i_{1}=0.794\angle 111^{\circ}\text{ A}$,  $i_{2}=0.790\angle-61.7^{\circ}\text{ A}$,  $i_{3}=0.229\angle 176^{\circ}\text{ A}$
Figure 11: An example in our dataset including a student’s solution on using linear algebra and complex operation to solve the target currents.
Example: Gemini-2.5-Pro transcribed student handwritten solution in Figure 6
Figure 12: Transcription generated by Gemini-2.5-Pro from the student handwritten solution shown in Figure 6. Expert-identified recognition errors are highlighted in red.
Example: Expert-rectified Gemini-2.5-Pro transcribed student handwritten solution in Figure 6
Figure 13: Expert-rectified version of the Gemini-2.5-Pro transcription from the student handwritten solution shown in Figure 6. Corrections corresponding to the original recognition errors are highlighted in green.
Example: GPT-5.1 transcribed student handwritten solution in Figure 7
Figure 14: Transcription generated by GPT-5.1 from the student handwritten solution shown in Figure 7. Expert-identified recognition errors are highlighted in red.
Example: Expert-rectified Gemini-2.5-Pro transcribed student handwritten solution in Figure 7
Figure 15: Expert-rectified transcription from Gemini-2.5-Pro for the student handwritten solution shown in Figure 7. The highlighted segments in green correspond to parts where GPT-5.1 (used here as the target model for diversity in few-shot learning) produced recognition errors, as shown in Figure 14. The LLM-based error detector is expected to produce a recognition error report similar to that shown in Figure 19 when comparing these transcriptions.
Prompt for student handwritten solution transcription
Figure 16: Prompt used to instruct Gemini-3-Pro-Preview, Gemini-2.5-Pro, GPT-5.1, and Claude-4.5-Sonnet to transcribe student handwritten solutions into Markdown format.
Prompt for student handwritten solution transcription (for Qwen3-VL series)
Figure 17: Prompt used to instruct Qwen3-VL-Plus and Qwen3-VL-8B-Thinking to transcribe student handwritten solutions into Markdown format.
One-shot prompt for the LLM detector to find recognition errors in the MLLM recognized text
Figure 18: Prompt used by the LLM-enabled recognition error detector to identify potential recognition errors in the target transcription (i.e., target_content) by comparing it with the expert-verified gold standard (i.e., label_content). To enhance detection performance, a one-shot learning approach is employed by providing an example that includes expert-annotated recognition errors and their corresponding rectifications.
One-shot example for LLM recognition error detector: errors and their rectified versions
Figure 19: Recognition errors and their corresponding rectifications for the sample shown in Figure 7, based on the comparison between the MLLM-generated transcription and the expert-verified gold standard. In this example, six item-level recognition errors were identified and corrected. The output format of the LLM-based recognition error detector follows this same structure.
One-shot example for LLM recognition error detector: explanations for capturing some errors
Figure 20: Explanations for each item-level recognition error presented in Figure 19. These justifications serve as reference examples to guide the LLM-based detector in identifying and interpreting recognition errors.
A rubric example for question P9.7-2 in EDU-CIRCUIT-HW
Figure 21: Grading rubric for question P9.7-2, outlining how to identify student mistakes and assign score deductions across five categories: “E”, “M”, “C”, “U”, and “NC”, as defined in Table 4.
Prompt (Round 1) for LLM grading (generate deducted scores if any)
Figure 22: First-round prompt used by the LLM grader to generate point deductions based on detected mistakes in the student’s solution.
Prompt (Round 2) for LLM grading (explaining the score deductions)
Figure 23: Prompt for the second round in which the LLM grader explains the reasons for score deductions and identifies their locations within the student’s transcribed Markdown solution.
Prompt for the LLM regrader to grade student solutions given the flagged recognition errors
Figure 24: Prompt used to instruct the LLM regrader (GPT-5.1) to evaluate student solutions while accounting for recognition errors flagged by the recognition error detector.
Prompt for heuristically identifying recognition errors from the target recognized text (Part I)
Figure 25: Prompt (Part I) for heuristically identifying potential recognition errors in transcriptions of student handwritten work from EDU-CIRCUIT-HW. The prompt includes a set of rules to help the detector exclude genuine student mistakes from flagged recognition errors. If the detector is uncertain about a particular case, it labels it as “uncertain”. During the regrading phase, all such cases are reviewed by a human expert (the doctoral teaching assistant in our case study).
Prompt for heuristically identifying recognition errors from the target recognized text (Part II)
Figure 26: Prompt (Part II) for heuristically identifying potential recognition errors in transcriptions of student handwritten work from EDU-CIRCUIT-HW. The markdown_content field contains the MLLM-generated transcription to be analyzed, while problem_statement and final_answer provide question-specific context (as introduced in Appendix B) to support accurate error detection.